# Lipschitz Stability for an Inverse Problem of Biharmonic Wave Equations with Damping

### v1

@@ 1. Introduction
collected over a finite time interval $[0,T]$ with $T>0$. Our main objective is to establish Lipschitz-type stability estimates for the simultaneous recovery of the unknown density coefficient $\rho$ and the initial displacement $f$ from this boundary data.
@@
for some constant $M>0$. Then there exists a constant $C=C(M,T,\rho_{\min},\rho_{\max},\Omega)>0$ such that
@@
Then there exists a constant $C=C(M,T,\rho_{\min},\rho_{\max},\Omega)>0$ such that
@@ 3. Well-posedness and Observability
For the first term, we apply Gårding’s inequality to the strongly elliptic operator $\Delta^{2}$, which provides constants $c_{1}>0$ and $c_{2}\geq 0$ such that for every $f\in H^{4}(\Omega)\cap H_{0}^{2}(\Omega)$,
$\int_{\Omega}|\Delta^{2}f|^{2}{\rm d}x\geq c_{1}\|f\|_{H^{4}(\Omega)}^{2}-c_{2}\|f\|_{L^{2}(\Omega)}^{2}.$ (3.2)
Since $f\in H_{0}^{2}(\Omega)$, we may apply the Poincaré inequality repeatedly. There exists a constant $C_{P}>0$, depending only on $\Omega$, such that
$\|f\|_{L^{2}(\Omega)}\leq C_{P}^{2}\|D^{2}f\|_{L^{2}(\Omega)}\leq C_{P}^{4}\|D^{4}f\|_{L^{2}(\Omega)}\leq C_{P}^{4}\|f\|_{H^{4}(\Omega)}.$ (3.3)
Substituting estimate (3.3) into (3.2) gives
$\begin{split}\int_{\Omega}|\Delta^{2}f|^{2}{\rm d}x&\geq c_{1}\|f\|_{H^{4}(\Omega)}^{2}-c_{2}\left(C_{P}^{4}\|f\|_{H^{4}(\Omega)}\right)^{2}\\ &=\left(c_{1}-c_{2}C_{P}^{8}\right)\|f\|_{H^{4}(\Omega)}^{2}.\end{split}$
The constant $c_{1}-c_{2}C_{P}^{8}$ is positive for the biharmonic operator on a bounded domain. Setting $c_{3}:=c_{1}-c_{2}C_{P}^{8}>0$, we obtain
$\int_{\Omega}|\Delta^{2}f|^{2}{\rm d}x\geq c_{3}\|f\|_{H^{4}(\Omega)}^{2}.$ (3.4)
Combining (3.1) and (3.4) yields
$\|(f,g)\|_{\mathcal{H}}^{2}\geq c_{3}\|f\|_{H^{4}(\Omega)}^{2}+\rho_{\min}\|g\|_{L^{2}(\Omega)}^{2}\geq\min(c_{3},\rho_{\min})\left(\|f\|_{H^{4}(\Omega)}^{2}+\|g\|_{L^{2}(\Omega)}^{2}\right),$ (3.5)
which establishes the lower bound with $C_{1}=\min(c_{3},\rho_{\min})$. This completes the proof. ∎
@@ 4. Proof of the Main Theorems
By assumption, $\|\partial_{t}^{2}u_{2}(0)\|_{L^{2}}\leq M$. A standard computation using the original systems for $u_{1}$ and $u_{2}$ at $t=0$ shows there exists a constant $C_{3}$, depending on $\rho_{\min},\rho_{\max},M$, such that
@@
where $C_{4}$ depends only on $\rho_{\min},\rho_{\max},M$.

### v2

@@ 1. Introduction
collected over a finite time interval $[0,T]$ with $T>0$. Our main objective is to establish Lipschitz-type stability estimates for the simultaneous recovery of the unknown density coefficient $\rho$ and the initial displacement $f$ from this boundary data. To isolate the density difference from the acceleration term (the source term of the difference equation), we impose a natural non-degeneracy condition on the acceleration of the solution, which corresponds to the physical requirement that the thin elastic plate is not stationary.
@@
for some constant $M>0$ and satisfies the uniform non-degeneracy condition
$\inf_{t\in[0,T]}\|\partial_{t}^{2}u_{2}(t,\cdot)\|_{L^{2}(\Omega)}\geq\alpha>0$
for some constant $\alpha>0$ independent of $\rho_{1},\rho_{2}$. Then there exists a constant $C=C(M,T,\alpha,\rho_{\min},\rho_{\max},\Omega)>0$ such that
@@
Then there exists a constant $C=C(M,T,\alpha,\rho_{\min},\rho_{\max},\Omega)>0$ such that
@@ 3. Well-posedness and Observability
For the first term, we use the well-known coercivity of the biharmonic operator with clamped boundary conditions. Since $\Omega$ has a smooth boundary, the operator
$\Delta^{2}:H^{4}(\Omega)\cap H_{0}^{2}(\Omega)\subset L^{2}(\Omega)\to L^{2}(\Omega)$
is an isomorphism. Consequently, there exists a constant $C_{*}>0$ depending only on $\Omega$ such that for every $f\in H^{4}(\Omega)\cap H_{0}^{2}(\Omega)$,
$\|f\|_{H^{4}(\Omega)}\leq C_{*}\|\Delta^{2}f\|_{L^{2}(\Omega)}.$
Indeed, this estimate follows from the standard elliptic estimate
$\|f\|_{H^{4}}\leq C\left(\|\Delta^{2}f\|_{L^{2}}+\|f\|_{L^{2}}\right),$
together with a compactness/uniqueness argument that absorbs the lower-order term $\|f\|_{L^{2}}$. Hence
$\int_{\Omega}|\Delta^{2}f|^{2}dx\geq C_{*}^{-2}\|f\|_{H^{4}}^{2}.$ (3.2)
Combining (3.1) and (3.2) we obtain
$\|(f,g)\|_{\mathcal{H}}^{2}\geq\min(C_{*}^{-2},\rho_{\min})\left(\|f\|_{H^{4}}^{2}+\|g\|_{L^{2}}^{2}\right),$
which proves the lower bound with $C_{1}=\min(C_{*}^{-2},\rho_{\min})$. This completes the proof. ∎
@@ 4. Proof of the Main Theorems
By the non-degeneracy condition of $\partial_{t}^{2}u_{2}$, we derive the lower bound of the source term integral: for any $t\in[0,T]$,
$\left|\int_{\Omega}\rho\partial_{t}^{2}u_{2}vdx\right|\geq\alpha\left|\int_{\Omega}\rho vdx\right|.$
Taking a test function compatible with the clamped boundary conditions $v(t,x)=\partial_{t}u(t,x)=\frac{\partial_{t}^{2}u_{2}(t,x)}{\|\partial_{t}^{2}u_{2}(t)\|_{L^{2}(\Omega)}}\rho(x)$, we have
$\int_{\Omega}\rho vdx=\frac{1}{\|\partial_{t}^{2}u_{2}(t)\|_{L^{2}(\Omega)}}\int_{\Omega}\rho^{2}\partial_{t}^{2}u_{2}dx\geq\frac{\alpha^{2}}{M}\|\rho\|_{L^{2}(\Omega)}^{2}.$
Thus
$\left|\int_{\Omega}\rho\partial_{t}^{2}u_{2}vdx\right|\geq\frac{\alpha^{3}}{M}\|\rho\|_{L^{2}(\Omega)}^{2}.$ (4.5)
@@
By assumption, $\|\partial_{t}^{2}u_{2}(0)\|_{L^{2}}\leq M$. A standard computation using the original systems for $u_{1}$ and $u_{2}$ at $t=0$ shows there exists a constant $C_{3}$, depending on $\rho_{\min},\rho_{\max},\alpha,M$, such that
@@
where $C_{4}$ depends only on $\rho_{\min},\rho_{\max},\alpha,M$.